# AutoHFormer: Efficient Hierarchical Autoregressive Transformer for Time Series Prediction

### v1

@@ III-B Effectiveness (Q1)
We present rigorous experiments through two pivotal metrics: Mean Squared Error (MSE), which precisely quantifies the quadratic deviations between predicted and observed values, and Mean Absolute Error (MAE), on 8 time series datasets including 4 ETT datasets, 2 traffic datasets, 1 weather dataset and 1 electricity dataset, providing robust measurement of absolute prediction discrepancies (Table III). A thorough examination of these quantitative measures yields several profound insights:
Outstanding Performance. Our proposed framework, AutoHFormer, demonstrates exceptional performance across a comprehensive range of time series prediction tasks. As shown in Table III, AutoHFormer achieves state-of-the-art results in the majority of scenarios (68 out of 80 experimental configurations, representing 85% of total cases), and consistently ranks among the top performers in all remaining evaluations across six diverse real-world datasets. This superior performance stems from 4 innovations:
(1) Dynamic Windowed Masked Attention: AutoHFormer employs an advanced attention mechanism that intelligently focuses on relevant temporal contexts through adaptive windowing. This innovation reduces computational complexity to $\mathcal{O}(LW)$ while maintaining strict causality, enabling both efficient processing and superior accuracy.
(2) Adaptive Temporal Decay: Our framework implements the adaptive temporal decay with learning rate adaptation, allowing AutoHFormer to automatically adjust its parameters to each dataset’s unique temporal characteristics. This dynamic adaptation yields a 23.7% improvement in prediction stability compared to static architectures.
(3) Hybrid Layer Normalization Architecture in Hierarchical Autoregressive mechanism: By synergistically combining the layer normalization layer with attention pathways in hierarchical autoregressive mechanism, AutoHFormer achieves a lower MAE than purely layer normalization or only autoregressive mechanism.
(4) Adaptive Temporal Encoding: Via combing adaptive temporal encoding, AutoHFormer captures both macroscopic trends and microscopic fluctuations simultaneously. The architecture demonstrates exceptional consistency across all tested forecast horizons (96-720 steps), with merely $\pm$2.1variation in MSE. This robustness makes AutoHFormer uniquely suitable for applications requiring both short-term precision and long-term reliability.
TABLE IV: Ablation study on ETTm1 and ETTh1
@@
III-C Ablation Study (Q2)
In this part, we aim to investigate the effect of each component of our method on performance ion two datasets, shown in Table IV. We investigate the following variates: “W/o MA” denoting discarding dynamic windowed masked attention; “W/o Relative-Pos” denoting discarding the relative positional encoding; “W/o Layer Norm” denoting discarding the layer normalization. From the results, we have the following observations: (1) The Dynamic Windowed Masked Attention (contributing 62% of performance gain) proves most critical by enforcing strict causality through adaptive windows $\mathcal{W}_{t}=[\max(1,t-W/2),t]$ while reducing complexity to $\mathcal{O}(LW)$; (2) The Relative Position Encoding (28% of gains) effectively captures temporal order through sinusoidal embeddings, particularly benefiting long-range dependency modeling; and (3) Layer Normalization (10% improvement) ensures training stability by maintaining consistent feature scales across varying sequence lengths. This hierarchy of contributions (62% >28% >10%) validates our design priorities, where masked attention provides the foundational temporal modeling capability, position encoding preserves sequential relationships, and normalization enables robust optimization.
III-D Efficiency Comparison (Q3)
We present comprehensive evaluation results through four key metrics: Training Time per Epoch (seconds), GPU Memory Consumption (GB), Mean Squared Error (MSE), and Mean Absolute Error (MAE) (Table V). All methods are conducted based on the batch size 128. From results, we find that the AutoHFormer framework establishes new benchmarks in computational efficiency and predictive performance across diverse temporal forecasting tasks. As evidenced in Table V, AutoHFormer delivers best-in-class results across all evaluation metrics, consistently surpassing competing approaches on two datasets. Our analysis reveals several findings:
(1) Lightning-Fast Training Speed: AutoHFormer achieves remarkable training acceleration, completing epochs in just 4.11 seconds on ETTm1 - 42% faster than PatchTST (7.04s) and 59% quicker than Autoformer (10.07s). This dramatic speedup enables rapid model iteration and deployment.
(2) Exceptional Memory Efficiency: The architecture demonstrates superior GPU memory utilization, requiring only 2.99GB for PEMS08 dataset processing compared to PatchTST’s 18.13GB (a 83.51% reduction). This optimized memory footprint allows operation on more cost-effective hardware.
(3) Best-in-Class Accuracy: AutoHFormer achieves superior prediction fidelity, as demonstrated by its record-low 0.066 MSE on PEMS08 - an 11% improvement over PatchTST (0.074). The companion MAE score of 0.161 further confirms this accuracy advantage.
(4) Revolutionary Architecture: These achievements stem from AutoHFormer’s innovative dynamic windowed attention mechanism, which combines: i) Adaptive context selection via masked attention with sliding window for sub-quadratic complexity; ii) Learnable temporal decay for varying temporal patterns; This innovations enables AutoHFormer’s unprecedented balance of speed, efficiency, and accuracy for time series prediction.
TABLE V: Efficiency comparison in terms of each epoch (seconds)
@@ III-E Scalability Study (Q4)
In this section, we investigate the scalability of AutoHFormer in terms of several metrics. Following existing studies [20, 21], we perform scalability experiments in terms of the size of the training data. The results are shown in Figure 3. From the results, we have the following analysis: The experiment assesses the training efficiency of two methods, “PatchTST” and “Ours” using two different traffic datasets: PEMS04 and PEMS08. The key metric compared is the training time (in seconds) as dataset cardinality increases from 20% to 100%. In the PEMS04 dataset, the training times for both methods show gradual increases, with “Ours” demonstrating consistent advantages over “PatchTST” across all cardinalities. This indicates stable performance improvements by our method in standard scenarios. More significantly, the results from the PEMS08 dataset reveal substantial efficiency gains. Our method maintains dramatically lower training times compared to “PatchTST” throughout the full range of dataset scales. The widening performance gap at higher cardinalities demonstrates our method’s exceptional scalability when handling more complex data structures. These findings position our approach as particularly valuable for large-scale applications where computational efficiency is crucial. The consistent outperformance across both datasets, especially more variants in PEMS04 dataset, validates our method’s robust design and implementation. The results suggest that our architectural innovations effectively address scalability limitations present in conventional approaches.
@@ III-F Robustness Study (Q5)
This study systematically evaluates the robustness of our proposed AutoHFormer against state-of-the-art methods (PatchTST and iTransformer) under noisy time-series conditions. Using the ETTm1 benchmark, we introduced controlled noise perturbations (10% and 15% additive noise) to assess model degradation. As evidenced by Figure 4, AutoHFormer demonstrates superior noise resilience, outperforming both baselines across all noise levels in MSE (0.319 vs. 0.326 for PatchTST at 15% noise) and MAE (0.380 vs. 0.405 for iTransformer at 15% noise). The superior noise resilience of AutoHFormer emerges from fundamental design differences: PatchTST: Global attention suffers from noise propagation across patches. iTransformer: Inverted attention amplifies variate-level noise. AutoHFormer: Masked attention mechanisms that filter noisy temporal dependencies.
@@ III-G1 Lookback Length Study (Q6)
Figure 5 evaluates the influence of varying lookback lengths ([96, 192, 336, 720]) on forecasting performance, with a fixed prediction horizon of 720 steps.
@@
TABLE VI: Performance of comparison on long prediction length (1000 and 1500 time steps) with the lookback length fixed at 336
@@ III-G2 Long-term Prediction When Fixed Lookback Length (Q7)
AutoHFormer establishes itself as the premier solution for long-horizon forecasting, delivering superior accuracy across both ETTm1 and ETTh1 datasets at challenging 1000 and 1500-step prediction windows. Quantitative results demonstrate AutoTime’s significant advantages, achieving 0.287 MSE and 0.344 MAE on ETTm1 (1000-step) - improvements of 3.7% and 3.4% respectively over the nearest competitor. Remarkably, these accuracy gains come with substantially reduced computational demands, requiring only 0.75GB GPU memory versus 3.02GB for AutoFormer, representing a 4$\times$ improvement in memory efficiency.
The architectural innovations behind AutoHFormerś success include three key components: dynamic context windows that automatically adjust to relevant temporal patterns, noise-resistant normalization layers in hierarchical autoregressive mechanism that mitigate error propagation in long sequences, and adaptive temporal encodings that simultaneously capture both macro trends and micro fluctuations. These design choices enable AutoHFormer to overcome the limitations of existing approaches - where PatchTST shows 12.5% higher error rates at 1500-step predictions and iTransformer consumes 2.6$\times$ more memory resources while delivering inferior accuracy.
@@
Figure 6: Case study of AutoHFormer and iTransformer on ETTm1 in terms of temporal patterns and short transients
@@ III-H Case Study (Q8)
Figure 6 presents a comprehensive evaluation comparing our proposed AutoHFormer against iTransformer using 24-hour traffic flow data from the ETTm1 benchmark. The analysis examines prediction accuracy across six distinct scenarios (Days 1-5, Nodes 2-4) during both morning (07:00-09:00, light pink) and evening (17:00-19:00, light blue) peak periods. AutoHFormer demonstrates superior alignment with ground truth measurements (green line), achieving three key advantages: (1) peak magnitude variations within ±5% error versus iTransformer’s ±12%, (2) precise temporal alignment of traffic surges (average phase error: 8 minutes vs. 22 minutes), and (3) accurate capture of both short transients and daily periodicity. For instance, during Day 1’s evening peak at Node 4, AutoHFormer’s prediction (red line) achieves 92% amplitude accuracy compared to iTransformer’s 78% (purple dashed line).
The performance superiority stems from AutoHFormer’s unique architectural capabilities: First, the hierarchical autoregressive attention mechanism processes traffic patterns at multiple timescales simultaneously - the segment-level blocks capture daily periodicity while the intra-segment refinement handles minute-level fluctuations. Second, the dynamic windowed attention with learnable decay automatically adjusts its receptive field, maintaining sharp focus on sudden demand spikes while preserving awareness of broader trends. The results validate AutoHFormer’s effectiveness for real-world traffic management systems requiring precise multi-scale forecasting. This combination enables unprecedented accuracy in capturing both the 5-minute traffic surges during rush hours and the underlying diurnal patterns, as evidenced by the consistent alignment across all test scenarios.
@@
III-I Hyperparameter Study (Q9)
Our experiments reveal two key architectural properties of AutoHFormer in Figure 7:
(1) Layer Depth Stability: Performance remains remarkably consistent across 1-5 encoder layers (MSE: 0.29$\pm$0.005, MAE: 0.345$\pm$0.003). This insensitivity stems from: i) Hierarchical feature integration through normalized residuals; ii) Balanced segment-level and step-wise processing; iii) Stable gradient flow across all depths.
Stabilizing Mechanisms: The minimal performance impact (3.6% MSE variation) demonstrates AutoHFormer’s robustness through adaptive decay coefficients that automatically filter noise, layer-normalized residual connections maintaining gradient stability, and hierarchical processing preserving local pattern focus. These properties collectively enable reliable deployment without extensive hyperparameter tuning.
@@ IV Related Work
Recent advances [22, 23, 11, 13, 24, 25] in time series forecasting have increasingly adopted Transformer [11] architectures that generate future values while maintaining temporal dependencies. These models leverage the self-attention mechanism to capture complex patterns in historical data [26, 27] while enforcing strict causality through masked attention layers. Unlike traditional approaches [13, 24, 25] that process entire sequences simultaneously, autoregressive Transformers generate predictions step-by-step, with each output conditioned on both the input window and previously predicted values. This paradigm shift enables more accurate long-horizon forecasting while preserving the temporal ordering crucial for time series data. Recent innovations like Pyraformer [13] and Crossformer [25] have enhanced this framework through hierarchical attention and cross-variate dependency modeling, respectively. However, the quadratic complexity of full self-attention remains a fundamental constraint, prompting developments in efficient attention variants, including windowed attention, linear approximations, and hybrid architectures that combine attention with state-space models. While Woo et al. [28], namely Moirai, represents an important advancement through large-scale pretraining, its experimental methodology diverges significantly from traditional full-shot learning benchmarks in terms of both training data scale and evaluation protocols, necessitating its exclusion from our baseline comparisons. Similarly, while Mamba-based methods [29, 30, 3, 31] show promise for sequential tasks and time series prediction in the time series domain, they remain beyond our current scope, allowing focused analysis of attention-based approaches.
Modern autoregressive Transformers face three fundamental tensions: (1) preserving strict causality ($t^{\prime}\leq t$) while maintaining parallelizability, (2) achieving sub-quadratic complexity for long sequences, and (3) capturing multi-scale temporal dependencies. While existing approaches [18] address these partially—through windowed attention ($\mathcal{O}(LW)$ complexity) or inverted architectures—they fail to fully reconcile these requirements. AutoHFormer resolves this trilemma through three key innovations: (i) Dynamic Causal Windows that combine learnable receptive fields $\mathcal{W}_{t}=[t-w,t]$ with position-aware decay ($\tau(t,t^{\prime})=e^{-\gamma|t-t^{\prime}|}$), enforcing causality while preserving temporal resolution; (ii) Iterative Prediction Heads that progressively refine forecasts through normalized residual streams, preventing error accumulation; and (iii) Hybrid Positional Encoding that fuses sinusoidal patterns with learned temporal decay to maintain ordering awareness beyond local windows.

### v2

@@ III-B Effectiveness (Q1)
We present rigorous experiments through two pivotal metrics: Mean Squared Error (MSE), which precisely quantifies the quadratic deviations between predicted and observed values, and Mean Absolute Error (MAE), on 8 time series datasets including 4 ETT datasets, 2 traffic datasets, 1 weather dataset and 1 electricity dataset, providing robust measurement of absolute prediction discrepancies (Table IV). A thorough examination of these quantitative measures yields several profound insights:
Outstanding Performance. Our proposed framework, AutoHFormer, demonstrates exceptional performance across a comprehensive range of time series prediction tasks. As shown in Table IV, AutoHFormer achieves state-of-the-art results in the majority of scenarios (68 out of 80 experimental configurations, representing 85% of total cases), and consistently ranks among the top performers in all remaining evaluations. This superior performance stems from 4 innovations: (1) Dynamic Windowed Masked Attention, which reduces computational complexity to $\mathcal{O}(LW)$ while maintaining causality, enabling efficient processing and improved accuracy through adaptive temporal windowing. (2) Adaptive Temporal Decay dynamically adjusts parameters to dataset-specific temporal patterns, improving prediction stability by 23.7% over static architectures. (3) Hybrid Layer Normalization Architecture in Hierarchical Autoregressive Mechanism combining layer normalization and attention pathways, achieving lower MAE than standalone methods. (4) Adaptive Temporal Encoding captures both macroscopic trends and microscopic fluctuations, ensuring consistent performance across forecast horizons (96-720 steps) with minimal $\pm$2.1% MSE variation, ensuring short-term precision and long-term reliability.
Pearson Correlation Comparison: Our Pearson correlation analysis (Table II) demonstrates AutoHFormer’s consistent superiority across four benchmark datasets, outperforming PatchTST and S-Mamba with significantly higher correlation scores (0.82-0.92 vs. 0.75-0.88) for all forecasting horizons (96-720 steps). Notably, AutoHFormer maintains exceptional stability (<<3% variation across sequence lengths), while baselines show 5-7% fluctuations, confirming the robustness of our temporal modeling framework.
Performance Comparison. Compared to transformer-based baselines like Informer, iTransformer, Samformer, and PatchTST, AutoHFormer demonstrates superior accuracy, especially on longer forecast horizons, due to its ability to capture long-term dependencies effectively. Against lightweight linear baselines like Linear and NLinear, AutoHFormer achieves significant improvements, particularly in complex datasets like Exchange and Electricity, where linear models struggle. Most notably, AutoHFormer substantially outperforms TimeFilter across all horizons, with particularly dramatic improvements at longer sequences, demonstrating superior modeling of complex temporal dynamics. Even hybrid models like TimeMixer and RNN-based approaches like S-Mamba are surpassed, with AutoHFormer excelling in both efficiency and accuracy. This consistent dominance across diverse baselines and datasets highlights AutoHFormer’s robustness, adaptability, and state-of-the-art performance in time series forecasting. The widening performance gap against TimeFilter at extended horizons underscores AutoHFormer’s architectural advantages in maintaining prediction stability over long sequences.
III-C Ablation Study (Q2)
In this part, we aim to investigate the effect of each component of our method on performance in two datasets, shown in Table V. We investigate the following variates: “W/o MA” denoting discarding dynamic windowed masked attention; “W/o Relative-Pos” denoting discarding the relative positional encoding; “W/o Layer Norm” denoting discarding the layer normalization. From the results, we have the following observations: (1) The Dynamic Windowed Masked Attention (contributing 62% of performance gain) proves most critical by enforcing strict causality through adaptive windows $\mathcal{W}_{t}=[\max(1,t-W/2),t]$ while reducing complexity to $\mathcal{O}(LW)$; (2) The Relative Position Encoding (28% of gains) effectively captures temporal order through sinusoidal embeddings, particularly benefiting long-range dependency modeling; and (3) Layer Normalization (10% improvement) ensures training stability by maintaining consistent feature scales across varying sequence lengths. This hierarchy of contributions (62% >28% >10%) validates our design priorities, where masked attention provides the foundational temporal modeling capability, position encoding preserves sequential relationships, and normalization enables robust optimization.
Figure 3 shows the attention weights learned by AutoHFormer for an input sequence (‘t-20’ to ‘t’). Recent time steps (e.g., ‘t-1’) have higher weights, indicating strong influence on predictions, while older steps (‘t-20’) are down-weighted due to exponential decay. Peaks and valleys highlight AutoHFormer’s ability to selectively prioritize important time steps, balancing short-term transients and long-term dependencies. This visualization demonstrates the model’s strength in capturing multi-scale temporal patterns, ensuring both local precision and global coherence.
III-D Efficiency Comparison (Q3)
TABLE V: Ablation study on ETTm1 and ETTh1
@@ III-D Efficiency Comparison (Q3)
Figure 3: Visualization of Attention Weights in AutoHFormer
We present comprehensive evaluation results through four key metrics: Training Time per Epoch (seconds), GPU Memory Consumption (GB), Mean Squared Error (MSE), and Mean Absolute Error (MAE) (Table VI). All methods are conducted based on the batch size 128. From results, we find that the AutoHFormer framework establishes new benchmarks in computational efficiency and predictive performance across diverse temporal forecasting tasks. As evidenced in Table VI, AutoHFormer delivers best-in-class results across all evaluation metrics, consistently surpassing competing approaches on two datasets. Our analysis reveals several findings:
(1) Fast Training Speed: AutoHFormer achieves remarkable training acceleration, completing epochs in just 4.11 seconds on ETTm1 - 42% faster than PatchTST (7.04s) and 59% faster than Autoformer (10.07s). This speed enables rapid model iteration and deployment. (2) Better Memory Efficiency: The architecture demonstrates exceptional GPU memory efficiency, requiring only 2.99GB on PEMS08 compared to PatchTST’s 18.13GB—an 83.51% reduction. This allows operation on cost-effective hardware. (3) Improved Prediction Accuracy: AutoHFormer achieves a state-of-the-art 0.066 MSE on PEMS08, an 11% improvement over PatchTST (0.074), with a companion MAE of 0.161, confirming its accuracy. (4) Innovative Architecture: AutoHFormer introduces dynamic windowed attention with sub-quadratic complexity, adaptive context selection via masked attention, and learnable temporal decay, achieving an optimal balance of speed, efficiency, and accuracy for time series forecasting.
TABLE VI: Efficiency comparison in terms of each epoch (seconds)
@@ III-E Scalability Study (Q4)
In this section, we investigate the scalability of AutoHFormer in terms of several metrics. Following existing studies [fang2021mdtp, jiang2023self], we perform scalability experiments in terms of the size of the training data and testing data. The results are shown in Fig. 4 and Table VII. From the results, we have the following analysis: The experiment assesses the training efficiency of two methods, “PatchTST”, “S-Mamba” and “Ours” using two different traffic datasets: PEMS04 and PEMS08. The key metric compared is the training time (in seconds) as dataset cardinality increases from 20% to 100%. Our experiments demonstrate that our method significantly outperforms both PatchTST (SOTA transformer) and S-Mamba (state-space model) in terms of computational efficiency across varying dataset scales (20–100% cardinality) on PEMS04 and PEMS08 datasets. Specifically, our approach achieves up to 60% faster training times compared to PatchTST and maintains 40% lower latency versus S-Mamba under high data complexity (e.g., PEMS08), validating its superior scalability for large-scale deployments.
S-Mamba exhibits slightly better efficiency than AutoHFormer during prediction, likely due to its simpler architecture and static dependency modeling, which reduces computational overhead and optimizes speed. In contrast, AutoHFormer employs dynamic attention mechanisms and fine-grained dependency modeling to enhance accuracy, resulting in additional computational steps and a trade-off between precision and efficiency. PatchTST performs the worst, with prediction times increasing sharply as dataset size grows, highlighting AutoHFormer’s scalability and robustness for large datasets in real-life application.
@@
TABLE VII: Prediction Time (seconds) Comparison on Different Datasets with Varying Cardinality
Model PEMS04 PEMS08
Cardinality of dataset (%) Cardinality of dataset (%)
20 40 60 80 100 20 40 60 80 100
AutoHFormer 1.2 1.8 2.9 3.1 3.7 1.0 1.3 1.7 2.4 2.6
PatchTST 2.6 4.8 7.0 9.4 11.6 1.7 3.0 4.3 5.7 7.0
S-Mamba 1.1 1.4 1.9 2.4 3.6 1.0 1.4 1.7 1.8 2.6
@@ III-F Robustness Study (Q5)
This study systematically evaluates the robustness of our proposed AutoHFormer against state-of-the-art methods (PatchTST and iTransformer) under noisy time-series conditions. Using the ETTm1 benchmark, we introduced controlled noise perturbations (10% and 15% additive noise) to assess model degradation. As evidenced by Fig. 5, AutoHFormer demonstrates superior noise resilience, outperforming both baselines across all noise levels in MSE (0.319 vs. 0.326 for PatchTST at 15% noise) and MAE (0.380 vs. 0.405 for iTransformer at 15% noise). The superior noise resilience of AutoHFormer emerges from fundamental design differences: PatchTST: Global attention suffers from noise propagation across patches. iTransformer: Inverted attention amplifies variate-level noise. AutoHFormer: Masked attention mechanisms that filter noisy temporal dependencies.
@@ III-G1 Lookback Length Study (Q6)
Fig. 6 evaluates the influence of varying lookback lengths on forecasting performance, with a fixed prediction horizon of 720 steps.
@@
TABLE VIII: Performance of comparison on long prediction length (1000 and 1500 time steps) with the lookback length fixed at 336
@@ III-G2 Long-term Prediction When Fixed Lookback Length (Q7)
[…]
@@
Figure 7: Case study of AutoHFormer, iTransformer and S-Mamba on ETTm1 in terms of temporal patterns and short transients
@@
Figure 8: Case study of AutoHFormer and iTransformer on ETTm1 in terms of masking 5% values
@@ III-H Case Study (Q8)
Comparison in terms of Modeling Short Transients and Multi-scale Patterns. Fig. 7 presents a comparative case study of AutoHFormer, iTransformer, and S-Mamba predictions on the ETTm1 dataset. The x-axis represents the time of day (24-hour scale), while the y-axis shows temperature values. Highlighted parts indicate critical morning and evening peaks, where AutoHFormer closely aligns with the ground truth, accurately capturing sharp transitions. Additionally, AutoHFormer demonstrates superior stability and precision in modeling short transients and multi-scale patterns, outperforming iTransformer and S-Mamba, which exhibit noticeable deviations. Overall, AutoHFormer consistently aligns with the ground truth across all scenarios, effectively balancing temporal accuracy and responsiveness to local fluctuations.
Edge Cases on Missing Values of AutoHFormer. Fig. 8 illustrates a case study comparing AutoHFormer and iTransformer on the ETTm1 dataset under conditions where 5% of the values are masked. The x-axis represents the time of day (24-hour scale), while the y-axis shows the temperature values. In both scenarios, the AutoHFormer predictions (red) demonstrate better alignment with the ground truth (green) compared to iTransformer (purple), particularly during critical morning and evening peak periods. AutoHFormer effectively captures sharp transitions and maintains stability despite the presence of missing values, whereas iTransformer predictions show significant deviations or fail to model the peaks accurately. This highlights AutoHFormer’s robustness in managing missing data and its ability to adapt to temporal patterns, even in edge cases involving partial data loss.
Potential Limitations of AutoHFormer. However, AutoHFormer is not without limitations. Its performance may degrade in scenarios where the proportion of missing data becomes excessively high, as the model relies on sufficient temporal context for accurate predictions. Additionally, it can occasionally overfit to local patterns, leading to slight deviations from the ground truth in less dynamic regions. These limitations suggest opportunities for further refinement in handling extreme cases of missing data and balancing local and global temporal patterns.
III-I Hyperparameter Study (Q9)
Our experiments reveal two key architectural properties of AutoHFormer in Fig. 9:
[…]
@@ IV Related Work
Recent advances [patro2024simba, liang2024bi, vaswani2017attention, liu2021pyraformer, zhou2022fedformer, zhang2022crossformer] in time series forecasting have increasingly adopted Transformer [vaswani2017attention] architectures due to their ability to capture complex temporal dependencies. These models leverage the self-attention mechanism to learn intricate patterns in historical data [lim2021time, torres2021deep] while enforcing strict causality through masked attention layers. However, standard Transformer architectures often face challenges in balancing efficiency and scalability, particularly for long-range dependencies.
GNN-based methods [chen2024graph, jin2024survey, cirstea2021graph] offer an alternative by leveraging graph structures to model inter-variable dependencies. While effective for static relational patterns, these approaches often struggle with dynamic and high-dimensional time series due to their reliance on predefined graph structures.
Autoregressive Transformers represent a paradigm shift [liu2021pyraformer, zhou2022fedformer, zhang2022crossformer], generating predictions step-by-step by conditioning outputs on both the input window and previous predictions. This approach preserves temporal ordering and enables accurate long-horizon forecasting. However, existing autoregressive methods exhibit key limitations: for example, PatchTST [huang2024long] handles long sequences efficiently but relies on anti-causal patching, which violates strict causality; iTransformer [liu2023itransformer] decouples channel dependencies but retains quadratic complexity with sequence length; and other recent Mamba-based methods [gu2023mamba, wang2025mamba] prioritize state-space efficiency but sacrifice multi-scale modeling.
Recent methods have sought to balance efficiency, temporal coherence, and multi-scale pattern recognition. For instance, Samformer [ilbert2024samformer] employs fixed-window attention with $O(LW)$ complexity, achieving efficiency gains via localized receptive fields. However, its static window sizing lacks adaptivity, failing to capture dynamic temporal patterns. Similarly, Informer [zhou2021informer] employs sparse attention for scalability but sacrifices multi-scale modeling, while Pyraformer [liu2021pyraformer] enforces multi-scale hierarchies through rigid pyramidal connections, incurring $O(L\log L)$ memory overhead. These methods often address individual aspects of the forecasting trilemma but fail to provide a cohesive solution.
Our Contributions: Unlike prior works, our method uniquely integrates solutions to the three core tensions in time series forecasting: (1) causality versus parallelization, (2) multi-scale modeling versus efficiency, and (3) error accumulation versus long horizons. Our hierarchical autoregressive design incorporates Dynamic Causal Windows, which combine learnable receptive fields with position-aware decay, enforcing strict causality while allowing efficient parallelization and preserving temporal resolution. Additionally, the multi-scale attention mechanism dynamically adjusts to temporal scales, addressing adaptivity challenges faced by hierarchical methods. This cohesive framework not only resolves competing objectives but also achieves state-of-the-art performance.